OpenStudy (anonymous):

Find the product of v=2i+6j and w=-5i-2j

OpenStudy (bahrom7893):

(2i+6j)*(-5i-2j)

OpenStudy (bahrom7893):

?

Parth (parthkohli):

Yup, vectors ^

OpenStudy (shamim):

do u mean dot product of ur given 2 vectors

OpenStudy (bahrom7893):

Dot or cross?

OpenStudy (shamim):

or u mean cross product of ur given 2 vectors

OpenStudy (anonymous):

its dot product sorry i left that out

Parth (parthkohli):

Aw.

OpenStudy (shamim):

ok

OpenStudy (shamim):

u know\[i.i=1\]

OpenStudy (shamim):

\[j.j=1,k.k=1\]

OpenStudy (shamim):

but\[i.j=0\]

OpenStudy (shamim):

\[j.k=0,i.k=0\]

Parth (parthkohli):

\[\vec v \cdot \vec w = (2\hat{i}+6\hat{j}) \cdot (-5\hat i - 2\hat j) = (2 \times -5) + (6 \times -2) = -10 - 12 = -22\]

OpenStudy (shamim):

do u need any explanation